OpenStudy (geekfromthefutur):

Help plz ive been stuck on this for an hour i give MEDAL A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 12(1.03)n Part A: When the scientist concluded his study, the height of the plant was approximately 16.13 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(n) represent? Part C: What is the average rate of change of the function f(n) from n = 3 to n = 10, and what does it represent?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (mrnood):

\[12\times1.03^{n}\] Is that the correct equation?

OpenStudy (geekfromthefutur):

Is that for part a b or c

OpenStudy (geekfromthefutur):

@MrNood

OpenStudy (mrnood):

It is a question - for you to clarify the form of the equation you have written..

OpenStudy (geekfromthefutur):

Yes that is the correct equation

OpenStudy (mrnood):

so you know that at the end, the height is 16.13 The domain means the values for n which are used. So f(n) = 16.13 = 12(1.03)^n so find a value for n which is close to this solution...

OpenStudy (geekfromthefutur):

Im still confused @MrNood

OpenStudy (mrnood):

solve this equation for n 16.13= 12(1.03)^n (start by dividing both sides by 12)

OpenStudy (geekfromthefutur):

So 12 divided 16.13 or 1.03

OpenStudy (mrnood):

OK - you need to be able to solve this equation 16.13= 12(1.03)^n SO first do what I said (start by dividing BOTH sides by 12) Do you know about logs? - you will need this next.

OpenStudy (geekfromthefutur):

Yes i do

OpenStudy (mrnood):

SO first do what I said (start by dividing BOTH sides by 12) Post the new equation here

OpenStudy (geekfromthefutur):

1.34=0.085

OpenStudy (kootra):

42!!! THE ANSWER TO THE UNIVERSE!!!!

OpenStudy (mrnood):

the Left side is correct - what did you do for the right side?

OpenStudy (geekfromthefutur):

I got 0.0858

OpenStudy (mrnood):

I said what did you DO, not what dis you get. we are trying to solve for n - so it is still an unknown \[\frac{ 12\times(1.03)^{n} }{ 12 }\]

OpenStudy (geekfromthefutur):

1.033

OpenStudy (geekfromthefutur):

I told you i suck at math

OpenStudy (mrnood):

you wouldn't say you suck at sports/English/driving/spelling - why be proud of 'sucking' at maths? The above equation is simple. Look at it again and write the answer:

OpenStudy (geekfromthefutur):

Well i multiplied 12 x 1,03 and i got 1345.66 and then i would divide that by 12 i think if im correct?

OpenStudy (mrnood):

Please take your time, and don't guess. I'll re-write the equation: \[\frac{ 12x }{12 }\] now see the similarity: \[\frac{ 12(1.03)^{n} }{ 12 }\] So simplify the above..

OpenStudy (anonymous):

How

OpenStudy (anonymous):

LOL, this was 7 months ago...